# Quantum Field Theory on Multifractal Spacetime: Varying Dimension and Ultraviolet Completeness

### v1

@@
Alessio Maiezza1,2∗, Juan Carlos Vasquez3†
@@
3Centro de Fisica Fundamental, Universidad de Los Andes, Merida, Venezuela
@@
Inspired by various quantum gravity approaches, we explore quantum field theory where spacetime exhibits scaling properties and dimensional reduction with changing energy scales, effectively behaving as a multifractal manifold. Working within canonical quantization, we demonstrate how to properly quantize fields in such multifractal spacetime. Our analysis reveals that a non-differentiable nature of spacetime is not merely compatible with quantum field theory but significantly enhances its mathematical foundation. Most notably, this approach guarantees the finiteness of the theory at all orders in perturbation theory and enables rigorous construction of the $S$-matrix in the interaction picture. The multifractal structure tames dominant, large-order divergence sources in the perturbative series and resolves the Landau pole problem through asymptotic safety, substantially improving the theory’s behavior in the deep ultraviolet regime. Our formulation preserves all established predictions of standard quantum field theory at low energies while offering novel physical behaviors at high energy scales.
@@ 1 Introduction
Consider a sheet of paper, modeled topologically as two-dimensional. When crumpled, however, it assumes a different effective fractal dimension. We can estimate this physically by measuring the crumpled paper’s density. Solid objects have densities scaling as $r^{-3}$ (where $r$ represents a characteristic length), yielding $d=3$ as a dimensionality estimate that coincides with the topological dimension. The crumpled paper, however, follows a different scaling relation, typically $r^{-d^{eff}}$, where $32<d^{eff}<$ 3. This simple example illustrates a crucial insight: the physical dimension may be neither trivial nor externally fixed. Rather, it requires a “physical estimator”, a concept central to this work, analogous to the density in our example.
@@
How should dimension differ from large-scale observations? Space-time could possess intrinsic fractal properties, causing fields to experience a different dimension than the topological one. The idea of a rough spacetime was first proposed by Wheeler who called it spacetime foam [3] – see also [4, 5]. We shall broadly refer to the roughness (non-differentiability) of spacetime as “fractal spacetime”. While string theory often suggests topological dimensions exceeding four (albeit warped), many quantum gravity approaches propose the opposite: dimensional reduction. Theories including Causal Dynamical Triangulations [6, 7, 8], Asymptotic Safety [9, 10], Causal Set Theory [11], and Loop Quantum Gravity [12] suggest that effective dimension flows continuously from four at large scales to two at short distances. Additional perspectives appear in [13, 14]. A model universe with variable dimension is presented in [15].
This article examines this phenomenon from the QFT perspective, broadening the seminal work of [16] on QFT formulation in fractal spacetime in many directions. The propagator, central to QFT, can serve as a dimensional estimator through running couplings. We investigate how canonical quantization should be modified to accommodate fractal space-time and dimensional reduction. Our findings reveal a straightforward generalization of canonical quantization that significantly enhances QFT consistency. First, the theory becomes finite at all orders in perturbation theory. Second, the behavior of the perturbative series is improved since the leading singularities in the Borel plane, namely, the ultraviolet (UV) renormalons, disappear. Third, also the UV Landau pole disappears so that any QFT model embedded in multifractal spacetime becomes asymptotically safe. We trace these improvements back to the Poincaré non-invariance of the vacuum, enabling robust $S$-matrix construction by circumventing limitations imposed by the Haag theorem [17]. The resulting theory agrees with the standard quantum field theory at low energies, while all the changes and improvements would appear at some high energy scale we denote $M$.
@@ 2 Fractality, smoothing and dimensional reduction
The convolution product $f*h$ gives
[…]
@@
For instance, think of the above function $f(x)$ as the distance between a horizontal line and a point in the 1D non-differentiable line given by the points belonging to the Koch curve. The function $f$ is continuous everywhere in $x$ but non-differentiable anywhere in $x$ [19]. Now assume that one is interested in finding the distance between any two points in the Koch curve. For zero resolution, such distance is not well-defined (infinite) when $l\rightarrow 0$, but it does exist for a finite, nonzero resolution $l$. After applying (2), the distance between two points in the Koch curve can be estimated and is given by
@@ 2.1 Introducing a scale in Euclidean Green equation: Dimensional Reduction
Let us now move to the case of a $\mathbb{R}^{d}$ non-smooth, Euclidean manifold. Similarly to the function $F$ in (2), the Green function has to depend on a scale that we call $M$, with the dimension of a mass-energy. The physical interpretation is that the space appears smooth at energy much below $M$, but it starts to show its fractal features at higher energy, comparable with $M$. This may be either an intrinsic scale of Nature or an effective energy emerging from a deeper theory than QFT, a theory that should include gravity at the quantum level  [^2] Fractal properties induced by quantum gravity effects are discussed in [20]..
Thus the standard Green equation,
@@
Therefore, $d^{eff}$ continuously flows from 4, at low energy, to $d^{eff}=2$ at $r\approx 2.2M^{-1}$, and it asymptotically reaches $d^{eff}=0$ at infinite energy. The interpretation is that $d^{eff}$ is an energy-dependent estimator of the non-constant dimension of the space. Specifically, the quantity $d^{eff}$ may be interpreted as an estimator of the “singularity spectrum” defining spacetime as a multi-fractal manifold [21, 22] – see [23, 24] for other applications to physical models. It is worth stressing that for energy close to $M$, the effective dimension becomes two, in agreement with many, if not all, suggestions from approaches to quantum gravity – see [25] for a comprehensive review. At energy much larger than $M$, the effective dimension vanishes. At energies much larger than the Planck energy, the effective dimension vanishes, which may not be immediately apparent. However, if we assume that $M$ is comparable to the Planck energy, we can speculate that the space-time dimensionality loses its meaning at energies where quantum fluctuations of space-time become dominant. This aligns with the analysis based on the spectral dimension estimator [26].
Before attempting the canonical quantization in this scenario, in the next section, let us summarize and comment on the picture we have introduced. First, the spacetime has topological dimension $d$ (e.g., $d=4$), thus one calculates in the standard way any algebra, vector, tensor, etc. usually appearing in QFT. Second, we conceptualize that quantum fields perceive an effective (varying) dimension $d^{eff}<d$, and this is defined by the behavior of the propagator for large momenta. The highlighted scenario shares some features with the one in [16], being one main difference that we consider and motivate a varying effective dimension. Another distinction is that we face the problem of rendering consistent canonical quantization in QFT, thereby dealing with well-known constraints.
@@ 3 Canonical quantization with varying effective dimension
Consider a real scalar field $\phi(x)$ and write the Feynman propagator,
@@
As in the usual QFT, it is convenient to write the classical field, before quantization, in the Fourier representation (fixing $d=4$)
$\phi(x,M)=\frac{1}{(2\pi)^{4}}\int\phi(p,M)e^{ipx}d^{4}p\,.$ (14)
Essentially, the standard quantization consists of replacing $\phi(p)$ with the ladder operators, $\phi(p)\rightarrow a(\vec{p}),a^{\dagger}(\vec{p})$ (from here on, we denote the 3-vector with $\vec{x}$ to distinguish it from $x$, denoting the 4-vector). In our case, we first recast $\phi(p,M)$ as
$\phi(p,M)=\sqrt{r(p,M)}\hat{\phi}(p)\,,$ (15)
where $r$ is a positive definite function, and $r\approx 1$ for $p\ll M$, reproducing standard QFT at low energy, but changing physics at deep UV. Then we have  [^3] To keep contact with [16], this expression can be regarded as a Stieltjes-Fourier transform of measure $d\mu_{H}:=\sqrt{r(p,M)}d^{4}p$.,
$\phi(x,M)=\frac{1}{(2\pi)^{4}}\int\sqrt{r(p,M)}\hat{\phi}(p)e^{ipx}d^{4}p\,.$ (16)
Next, we promote $\hat{\phi}(p)$ as an operator, finally writing it in terms of the ladder operator. Specifically, we have,
@@
One implements the standard canonical commutation relations (CCR) but on the field $\hat{\phi}$ in (15),
$[a(\vec{p}),a^{\dagger}(\vec{p^{\prime}})]=\delta(\vec{p}-\vec{p^{\prime}})\hskip 10.00002pt\Leftrightarrow\hskip 10.00002pt[\hat{\phi}(t,\vec{x}),\hat{\pi}(t,\vec{y})]=i\delta(\vec{x}-\vec{y})\,,$ (19)
where $\hat{\pi}=\dot{\hat{\phi}}$ is the conjugate variable of $\hat{\phi}$, and the square brackets denote the commutator.
One may also define,
$a(\vec{p},M)=\sqrt{r(\vec{p},M)}a(\vec{p})\,,$ (20)
such that CCR can be rewritten as,
$[a(\vec{p},M),a^{\dagger}(\vec{p^{\prime}},M)]=r(\vec{p},M)\delta(\vec{p}-\vec{p^{\prime}})=r(p,M)\delta(\vec{p}-\vec{p^{\prime}})\,,$ (21)
@@
[…]
@@
The generalization of the quantization, in the presence of dimensional reduction, of vector and fermion fields is straightforward. It is sufficient to decompose the fields as in (15) and then write covariant CCR, for vector fields, and anti-commutation relations (ACR) for fermions. The Reader might worry about the quantization of vector fields, then of gauge bosons, in the presence of a dimensional scale ($M$) since this may resemble a cutoff, thus inconsistent with gauge invariance. However, this is not the case. A sharp cutoff rules out the high momentum modes breaking the gauge invariance, conversely, the Fourier transforms of the fields, e.g., (17), ranges up to infinite energy.
@@ 3.1 Propagator
which has to be written in terms of 4-vectors. We achieve this in the usual way but with some attention to the function $r$. We call for shortness $s:=x-y$ and consider the integral,
[…]
@@
It is important to elaborate on the implications of (18) and (28). By comparing (8) and (29) (regardless that the latter is in Minkowski space) , one sees that $r$ is an even function in momentum. This will turn out to be problematic for the consistency of the theory. The function $r$ must be odd, so we can modify it as
@@
where $v$ is a constant 4-vector necessary to build $r$ as a function not only of $p^{2}$, and $V=\exp[v^{2}]$ is a normalization constant.
@@
However, the insight on dimensional reduction from (8) remains unchanged since it is only based on arguments of convergence or divergence of the propagator, dominated by the $p^{2}$ contribution, which reflects on the loop behavior or the couplings running. Notice also that $r$, in the form of (30), does not trivialize on-shell yielding effects even for tree-level processes.

### v2

@@
Alessio Maiezza1,2∗, Juan Carlos Vasquez3,4†
@@
3SISSA, Via Bonomea 265, I-34136 Trieste, Italy,
4Centro de Fisica Fundamental, Universidad de Los Andes, Merida, Venezuela.
@@
Inspired by various quantum gravity approaches, we explore quantum field theory where spacetime exhibits scaling properties and dimensional reduction with changing energy scales, effectively behaving as a multifractal manifold. Working within canonical quantization, we demonstrate how to properly quantize fields in such a multifractal spacetime. Our analysis reveals that a non-differentiable nature of spacetime is not merely compatible with quantum field theory but significantly enhances its mathematical foundation. Most notably, this approach ensures perturbative UV finiteness and improved behavior of the series expansion and enables rigorous construction of the $S$-matrix in the interaction picture by breaking vacuum translational invariance. The multifractal structure tames dominant, large-order divergence sources in the perturbative series and resolves the Landau pole problem through asymptotic safety, substantially improving the theory’s behavior in the deep ultraviolet regime. Our formulation preserves all established predictions of standard quantum field theory at low energies while offering novel physical behaviors at high energy scales.
@@ 1 Introduction
Consider a sheet of paper, modeled topologically as two-dimensional. When crumpled, however, it assumes a different effective fractal dimension. We can estimate this physically by measuring the crumpled paper’s density. Solid objects have densities scaling as $r^{-3}$ (where $r$ represents a characteristic length), yielding $d=3$ as a dimensionality estimate that coincides with the topological dimension. The crumpled paper, however, follows a different scaling relation, typically $r^{-d^{eff}}$, where $2<d^{eff}<3$. This simple example illustrates a crucial insight: the physical dimension may be neither trivial nor externally fixed. Rather, it requires a “physical estimator”, a concept central to this work, analogous to the density in our example.
@@
How should dimension differ from large-scale observations? Space-time could possess intrinsic fractal properties, causing fields to experience a different dimension than the topological one. The idea of a rough spacetime was first proposed by Wheeler, who called it spacetime foam [3] – see also [4, 5]. We shall broadly refer to the roughness (non-differentiability) of spacetime as “fractal spacetime”. While string theory often suggests topological dimensions exceeding four  [^2] Dimension reduction can also be formulated within string theory [6] and multibrane model [7]., many quantum gravity approaches propose the opposite: dimensional reduction. Theories including Causal Dynamical Triangulations [8, 9, 10], Asymptotic Safety [11, 12], Causal Set Theory [13], and Loop Quantum Gravity [14] suggest that effective dimension flows continuously from four at large scales to two at short distances. Additional perspectives appear in [15, 16]. A model universe with variable dimension is presented in [17]. For clarity, let us stress that the term ”(multi)fractal spacetime” in this work refers to an effective property derived from QFT propagator scaling, discussed below, and not to multifractional geometries [18, 19, 20, 21, 22, 23, 24]. We emphasize that we work with a flat, fractal space time, motivated from quantum gravity arguments, as discussed above.
This article examines this phenomenon from the QFT perspective, broadening the seminal work of [25] on QFT formulation in fractal spacetime in many directions. The propagator, central to QFT, can serve as a dimensional estimator through running couplings. We investigate how canonical quantization should be modified to accommodate fractal space-time and dimensional reduction. Our findings reveal a straightforward generalization of canonical quantization that significantly enhances QFT consistency. First, the theory becomes finite at all orders in perturbation theory. Second, the behavior of the perturbative series is improved since the leading singularities in the Borel plane, namely, the ultraviolet (UV) renormalons, disappear. Third, the UV Landau pole also disappears so that any QFT model embedded in multifractal spacetime becomes asymptotically safe. Finally, a remarkable feature is the Poincaré non-invariance of the vacuum, enabling robust $S$-matrix construction by circumventing limitations imposed by the Haag theorem [26].
While form factors, damping UV behavior, have been extensively studied in nonlocal quantum field theory and quantum gravity, motivated by string field theory and the need to control ultraviolet divergences while preserving unitarity and ghost-freeness [27, 28, 29, 30, 31, 32, 33, 34, 35, 36], here we modify the propagator, via a function (form factor) that would break the translational invariance of the vacuum, on which Haag’s theorem relies. Hence, enabling a consistent perturbative framework beyond what standard nonlocal models achieve.
@@ 2 Fractality, smoothing and dimensional reduction
The convolution product fhf*h gives
$F(x,l)=\int_{0}^{\infty}f(y)\,h(x-y)\,dy\,$, (2)
@@
For instance, think of the above function $f(x)$ as the distance between a horizontal line and a point in the 1D non-differentiable line given by the points belonging to the Koch curve. The function $f$ is continuous everywhere in $x$ but non-differentiable anywhere in $x$ [38]. Now, assume that one is interested in finding the distance between any two points in the Koch curve. For zero resolution, such a distance is not well-defined (infinite) when $l\rightarrow 0$, but it does exist for a finite, nonzero resolution $l$. After applying (2), the distance between two points in the Koch curve can be estimated and is given by
@@ 2.1 Introducing a scale in Euclidean Green equation: Dimensional Reduction
Let us now move to the case of a $\mathbb{R}^{d}$ non-smooth, Euclidean manifold. Similarly to the function $F$ in (2), the Green function has to depend on a scale that we call $M$, with the dimension of a mass-energy. The physical interpretation is that the space appears smooth at energy much below $M$, but it starts to show its fractal features at higher energy, comparable with $M$. This may be either an intrinsic scale of Nature or an effective energy emerging from a deeper theory than QFT, a theory that should include gravity at the quantum level  [^3] Fractal properties induced by quantum gravity effects are discussed in [39].. Our approach provides an operational definition of an effective dimension directly from the QFT propagator, distinct from the usual spectral dimension.
The standard Green for the Klein-Gordon equation,
@@
The replacement of $\delta^{(d)}(x-y)$ with a Gaussian kernel is consistent with the Schwinger proper-time representation and heat-kernel techniques widely used in QFT [40, 41, 42]. Form factors in nonlocal theories implement similar exponential suppression to achieve UV control without violating gauge invariance. In the framework discussed in this work, the Gaussian kernel models a resolution scale $M$, signaling the onset of multifractal features rather than imposing a hard cutoff in the theory.
@@
Therefore, $d^{eff}$ continuously flows from 4, at low energy, to $d^{eff}=2$ at $r\approx 2.2M^{-1}$, and it asymptotically reaches $d^{eff}=0$ at infinite energy. This dimensional reduction from 4 to 2 is consistent with findings from various quantum gravity approaches. The spectral dimension, defined via heat-kernel diffusion on quantum spacetimes [43, 44, 45], provides a measure of effective dimensionality. Our operational definition, based on the propagator scaling, differs from but complements the spectral dimension approach. The interpretation is that $d^{eff}$ is an energy-dependent estimator of the non-constant dimension of the space. Specifically, the quantity $d^{eff}$ may be interpreted as an estimator of the “singularity spectrum” defining spacetime as a multi-fractal manifold [46, 47] – see [48, 49] for other applications to physical models. It is worth stressing that for energy close to $M$, the effective dimension becomes two, in agreement with many, if not all, suggestions from approaches to quantum gravity – see [50] for a comprehensive review. At energy much larger than $M$, the effective dimension vanishes. At energies much larger than the Planck energy, the effective dimension vanishes, which may not be immediately apparent. However, if we assume that $M$ is comparable to the Planck energy, we can speculate that the space-time dimensionality loses its meaning at energies where quantum fluctuations of space-time become dominant. This aligns with the analysis based on the spectral dimension estimator [51].
Before attempting the canonical quantization in this scenario, in the next section, let us summarize and comment on the picture we have introduced. First, the spacetime has topological dimension $d$ (e.g., $d=4$), thus one calculates in the standard way any algebra, vector, tensor, etc., usually appearing in QFT. Second, we conceptualize that quantum fields perceive an effective (varying) dimension $d^{eff}<d$, and this is defined by the behavior of the propagator for large momenta. The highlighted scenario shares some features with the one in [25], being one main difference that we consider and motivate a varying effective dimension. Another distinction is that we face the problem of rendering consistent canonical quantization in QFT, thereby dealing with well-known constraints.
It is worth noting that nonlocal QFT and gravity models with damping functions have been developed in depth, addressing unitarity, super-renormalizability, and in some cases finiteness [30, 52, 53]. As we will see in the rest of this paper, our proposal differs conceptually by introducing a vacuum that is not translationally invariant, requiring a modification of the above form factor. In this way, we circumvent Haag’s theorem and establish a perturbative framework that is both UV finite and free of renormalon ambiguities.
@@ 3 Canonical quantization with varying effective dimension
Consider a real scalar field $\phi(x)$ with action
$S=\int d^{4}x\,\left(\frac{1}{2}(\partial_{\mu}\phi(x))^{2}-\frac{1}{2}m^{2}\phi(x)^{2}-\frac{\lambda}{4!}\phi(x)^{4}\right)\,$, (13)
and write the free Feynman propagator,
@@
As in the standard QFT, it is convenient to write the classical field, before quantization, in the Fourier representation (fixing $d=4$)
$\phi(x,M)=\frac{1}{(2\pi)^{4}}\int\phi(p,M)e^{ipx}d^{4}p\,$, (15)
which satisfies the Klein-Gordon equation, which is in momentum space
$(p^{2}-m^{2})\phi(p,M)=0\,.$ (16)
The solution is of the form,
$\phi(p,M)=2\pi\,\delta(p^{2}-m^{2})\,a(p,M)\,$, (17)
where $\delta$ is the Dirac’s delta function, and $a(p,M)$ is an arbitrary, non-singular function of the momentum and the parameter $M$, which we define as
$a(p,M):=a(p)\,\sqrt{r(p,M)}\,$, (18)
Then, the $M$-dependent field can be written as
$\phi(p,M)=\phi(p)\,\sqrt{r(p,M)}\,.$ (19)
Notice that this approach is radically different than the approach adopted in [49, 54], where the classical equations of motion are modified by a non-propagating field.
Essentially, the standard quantization consists in promoting $a(p,M)$ from a function to an operator, so defining the ladder operators with a dependence on the scale $M$.
The function $r$ is positive defined, and it is such that $r\approx 1$ for $p\ll M$, reproducing standard QFT at low energy, but changing physics at deep UV. Then we have  [^4] To keep contact with [25], this expression can be regarded as a Stieltjes-Fourier transform of measure $d\mu_{H}:=\sqrt{r(p,M)}d^{4}p$.,
$\phi(x,M)=\frac{1}{(2\pi)^{4}}\int\sqrt{r(p,M)}\phi(p)e^{ipx}d^{4}p\,$, (20)
which can be written explicitly as a non-local operator as:
$\phi(x,M)=\int d^{4}y\,R(x-y,M)\,\phi(y)\,$, (21)
where $R$ is the Fourier transform of $\sqrt{r(p,M)}$ given by
$R(x-y,M)=\int\frac{d^{4}p}{(2\pi)^{4}}\,e^{ip\cdot(x-y)}\,\sqrt{r(p,M)}\,.$ (22)
Next, we promote $\phi(p,M)$ (and $a(p,M)$) as quantum operators, finally writing it in terms of the ladder operator we have,
@@
One implements the standard canonical commutation relations (CCR) but on the field $\phi(p)$ in (19),
$a(\vec{p}),a^{\dagger}(\vec{p^{\prime}})]=(2\pi)^{3}\,\delta(\vec{p}-\vec{p^{\prime}})\hskip 10.00002pt\Leftrightarrow\hskip 10.00002pt[\phi(t,\vec{x}),\pi(t,\vec{y})]=i\delta(\vec{x}-\vec{y})\,$, (25)
where $\pi=\dot{\phi}$ is the conjugate variable of $\hat{\phi}$, and the square brackets denote the commutator. The CCR can be rewritten as,
$a(\vec{p},M),a^{\dagger}(\vec{p^{\prime}},M)]=(2\pi)^{3}\,r(\vec{p},M)\delta(\vec{p}-\vec{p^{\prime}})=(2\pi)^{3}\,r(p,M)\delta(\vec{p}-\vec{p^{\prime}})\,$, (26)
@@
$a^{\dagger}(\vec{p},M)|0\rangle=\sqrt{r(\vec{p},M)}|1_{\vec{p}}\rangle\hskip 20.00003pta(\vec{p},M)|0\rangle=0\,.$ (27)
@@
The generalization of the quantization, in the presence of dimensional reduction, of vector and fermion fields is straightforward. It is sufficient to decompose the fields as in (19) and then write covariant CCR, for vector fields, and anti-commutation relations (ACR) for fermions. The Reader might worry about the quantization of vector fields, then of gauge bosons, in the presence of a dimensional scale ($M$) since this may resemble a cutoff, thus inconsistent with gauge invariance. However, this is not the case. A sharp cutoff rules out the high momentum modes breaking the gauge invariance; conversely, the Fourier transforms of the fields, e.g., (23), range up to infinite energy.
@@ 3.1 Propagator
which has to be written in terms of 4-vectors. We achieve this in the usual way, but with some attention to the function $r$. We call for shortness s:=xys:=x-y and consider the integral,
$\frac{1}{(2\pi)^{4}}\int d^{3}\vec{p}e^{i\vec{p}\,\vec{s}}\int dp_{0}\frac{ir(p_{0},\vec{p},M)e^{-is_{0}\,p_{0}}}{p^{2}-m^{2}}\,.$ (31)
@@
It is important to elaborate on the implications of (24) and (33). By comparing (8) and (34) (regardless that the latter is in Minkowski space) , one sees that $r$ is an even function in momentum. This will turn out to be problematic for the consistency of the theory. The function $r$ must be asymmetric for $p\to-p$ (we shall refer to it simply as odd), so we can modify it as
@@
where $v$ is a constant 4-vector necessary to build $r$ as a function not only of $p^{2}$, and $V=\exp[v^{2}]$ is a normalization constant. This amounts to a redefinition of the form factor in momentum space in (7), namely, shifting $p\to p-M\,v$.
@@
However, the insight on dimensional reduction from (8) remains unchanged since it is only based on arguments of convergence or divergence of the propagator, dominated by the $p^{2}$ contribution, which reflects on the loop behavior or the couplings running. Notice also that $r$, in the form of (35), does not trivialize on-shell, yielding effects even for tree-level processes.